# Exploring AI-based System Design for Pixel-level Protected Health Information Detection in Medical Images

### v3

@@ Abstract
De-identification of medical images is a critical step to ensure privacy during data sharing in research and clinical settings. The initial step in this process involves detecting Protected Health Information (PHI), which can be found in image metadata or imprinted within image pixels. Despite the importance of such systems, there has been limited evaluation of existing AI-based solutions, creating barriers to the development of reliable and robust tools. In this study, we present an AI-based pipeline for PHI detection, comprising three key components: text detection, text extraction, and text analysis. We benchmark three models—YOLOv11, EasyOCR, and GPT-4o— across different setups corresponding to these components, evaluating the performance based precision, recall, F1 score, and accuracy. All setups demonstrate excellent PHI detection, with all metrics exceeding 0.9. The combination of YOLOv11 for text localization and GPT-4o for extraction and analysis yields the best results. However, this setup incurs higher costs due to GPT-4o’s token generation. Conversely, an end-to-end pipeline that relies solely on GPT-4o shows lower performance but highlights the potential of multi-modal models for complex tasks. We recommend fine-tuning a dedicated object detection model and utilizing built-in OCR tools to achieve optimal performance and cost-effectiveness. Additionally, leveraging language models such as GPT-4o can facilitate thorough and flexible analysis of text content.
@@ 1 Introduction
Protected Health Information (PHI) is health-related data that can identify an individual, often found in medical records, images, or any documents and data elements generated during participation in healthcare services. Before sharing data, e.g., for research purposes, it is crucial to de-identify or pseudonymize it to minimize the risk of re-identification. This de-identification step is mandatory for studies to comply with the Health Insurance Portability and Accountability Act (HIPAA) [1] and the General Data Protection Regulation (GDPR) [2], and it is the responsibility of the data provider. De-identification is a complex process that involves many data types and formats, typically including a detection step followed by the redaction of the identified PHI. Medical images are a rich source of PHI, which can be embedded as metadata (e.g., in DICOM headers) or burned-in text within pixel data. The automatic detection of pixel-level PHI is challenging due to the diversity in imprint content, locations, font sizes, color schemes, annotations, and machine overlays. As the overall volume of medical data increases, so does the risk of incomplete PHI redaction, potentially leading to data breaches, which imply significant negative consequences for affected individuals and the healthcare sector, as well as high costs for responsible entities [3]. The importance of reliable PHI detection at scale and automated image quality control in general is strongly emphasized by the latest medical image data platform development program launched by the Advanced Research Projects Agency for Health (ARPA-H) [4].
@@
Previous works have employed Optical Character Recognition (OCR) technology to recognize text in medical images, leading to two primary approaches for handling burned-in text. The first approach removes all burned-in texts, regardless of content [5]. While this method effectively eliminates any potential PHI, it also discards valuable information, such as measurements and markers. The second approach applies heuristic rules to redact specific PHI imprints [6]. However, this method is tailored to particular representations of imprints in images and can struggle to generalize to new types of PHI or different representations. In addition, fine-grained classification of PHI is often preferred to coarse classification, which categorizes information simply as names or dates. For instance, not all patient identifiers or dates need redaction. Identifiers linked to the hospital system pose a high risk of identification and should indeed be redacted, while those used in clinical trials are valid for study analysis. Similarly, although all dates are classified as PHI under HIPAA guidelines, certain dates, such as those documenting patient visits, are critical for clinical trials. As the volume and diversity of medical data continue to grow, advanced automatic solutions for PHI detection are essential, particularly for managing large datasets collected retrospectively. These solutions are expected to handle edge cases, which can involve various signaling words and phrases. Moreover, research sites require flexibility to determine which entities to include or exclude as PHI based on specific study setup. Therefore, a robust PHI detection tool must incorporate mechanisms that enable easy specification and adjustment of rules to meet the needs of different studies.
@@
Our study focuses on AI-based approaches for automatically detecting PHI as burned-in pixels. Currently, there are a few solutions that utilize vision and language models from MD.ai [7], Google [8], John Snow Labs [9], and Glendor [10]. These tools commonly integrate two key components: a vision component that detects and converts burned-in texts into machine-encoded strings, and a language component that identifies any PHI content in the extracted texts [11]. To achieve this, OCR models are utilized for text recognition, while name-entity recognition (NER) models are employed to classify the PHI content. Recently, large language models (LLMs) have become the preferred candidates for the language component due to their flexible configurability and ability to understand complex language context in NER tasks, which has proven helpful in detecting PHI in biomedical text [11]. Despite some general remarks concerning the deployment of these tools [12, 13], particularly on the non-deterministic behavior and cost of LLMs, there is a lack of comprehensive studies dissecting the setup of vision and language models within a PHI detection pipeline. In addition, as multimodal LLMs with vision capabilities are being released today [14], the question remains whether they can be used to recognize PHI through end-to-end processing of images. To address the above concerns, we make the following contributions in our study:
@@
  • By leveraging publicly available medical imaging data and the HIPAA guidelines, we create two evaluation datasets comprising 1000 and 200 images, respectively. These datasets encompass various radiology modalities, including X-Ray, CT, PET, MRI, bone scintigraphy, mammography, and ultrasound. The images are synthesized with pixel-level PHI to support development and benchmarking within the research community.
@@
  • Based on the quantitative results, we provide insights into the optimal pipeline setup, considering performance, latency, and inference cost in terms of the number of generated tokens. Specifically, we explore the full capabilities of GPT-4o for both vision and language tasks, and analyze the trade-offs involved in utilizing it partially or entirely in the PHI detection pipeline.
@@ 2 Datasets
For the study at hand, we generate and utilize three distinct datasets: RadPHI-train as our training dataset, and two evaluation datasets, RadPHI-test and MIDI. RadPHI-train and RadPHI-test are created by overlaying synthetically generated imprints on curated medical images across various modalities. MIDI is created by overlaying information stored in DICOM tags, which may contain PHI, utilizing a standard DICOM viewer. The RadPHI-train is used to develop our text detection model, while the RadPHI-test is employed to evaluate different configurations of the PHI detection pipeline. Additionally, the MIDI dataset is reserved as a hold-out test set to assess the optimal configuration identified through the RadPHI-test evaluation. In the upcoming sections, we will elaborate on the generation process for each of these datasets.
@@ 2.1 PHI definition
For data simulation and evaluation, we define a list of PHI and non-PHI items based on the HIPAA guidelines [1] as detailed in Table 1. The eight PHI categories include date, general identifier, patient name, address, age, gender, phone number, and email. While the HIPAA guidelines recommend that only ages over 89 are considered PHI, we classify any age as PHI. This is to avoid re-identification, as age groups can exhibit skewed distribution depending on the study, making participants identifiable through age data. Additionally, we group all identification numbers into a group of identifiers, including patient ID, insurance number, social security number, and any number or series that can be used to identify an individual. The eight non-PHI categories include height, weight, examination type, hospital, marker, scanner, diagnosis, and imaging personnel.
@@
Date All elements of dates related to an individual, including birth date, admission date, discharge date, etc. PHI DOB 01-01-2023
Identifier All identifiers related to an individual such as patient ID, insurance number, medical record number, etc. PHI Patient ID: 0000.0001
@@
Age The patient’s age at the time of the scan. PHI Age: 60
Gender The gender of an individual, e.g., male, female, or diverse, or the abbreviations such as m/f/d PHI [M]
@@
Table 1: Imprint categories and examples. There are 16 categories, eight of which are classified as PHI.
@@ 2.3 RadPHI-train
The main use of the RadPHI-train dataset is to enhance the robustness of the text localization module (see Section 3.1) against diverse radiological image backgrounds, varying numbers of imprints, and different imprint representations. This dataset is generated based on images from TotalSegmentator, BS-80K, and ChestX-ray8, applying the preprocessing techniques outlined in Section 2.2. Additionally, we simulate imprints on synthetic backgrounds, incorporating plain grayscale colors along with or without multiple boxes in various shades. The text localization module is trained to manage diverse representations of the imprint rather than focusing on its text content. To achieve this, we continuously simulate the imprints using different fonts, sizes, colors, and locations. These imprints may overlap with each other or with anatomical structures. Although such overlaps are not typically seen in real-world data, we intentionally include them as challenging cases for the model to adapt to. A sample from this dataset is shown in Figure C1 (left) in the Supplementary Material. In total, we generate 6,000 images, which are then divided into training and validation sets using an 80%-20% split.
@@ 2.4 RadPHI-test
Figure 5: Distribution of PHI and non-PHI imprints in RadPHI-test and MIDI datasets. In RadPHI-test, the number of imprints is limited to eight. In contrast, MIDI has significantly more imprints, correlating to the various DICOM tags randomly displayed on the viewer.
@@ 2.5 MIDI
The MIDI dataset is designed to be a challenging hold-out dataset with realistic imprint visualization and placement via a modern DICOM viewer rendering. The image data is curated from the validation set of the 2024 Medical Image De-Identification Benchmark (MIDI-B) challenge [19]. This dataset originally consists of 228 DICOM images across multiple modalities, each containing synthetic PHI content embedded at both the DICOM header and pixel level. We utilize a DICOM viewer, specifically MD.ai [20], to overlay the DICOM tags onto the images. We randomly sample DICOM tags to ensure that the generated imprints represent all possible PHI categories, similar to the RadPHI-test dataset. After applying the overlays, we export the images from the viewer. The resulting images may include not only the DICOM tag overlays but also burn-ins by the challenge organizers. The final version of the dataset comprises 200 images categorized into seven PHI types. The category Email is omitted because emails are neither burned in nor inserted into the header of the original dataset. In contrast to the RadPHI-test dataset, each image can have multiple imprints of the same PHI category. For instance, as illustrated in Figure 6, dates may appear in various contexts, such as the study date, image series, or even the patient comments. We annotate the images at the level of whether any PHI category is present; thus, even if there are numerous imprints of the same category, they are counted as a single instance in our annotation. Figure 4 displays the distribution of PHI categories at the image level. Dates and patient names are among the most frequently occurring PHI elements in the dataset. While we intentionally limit the number of imprints to eight in RadPHI-test, the number of imprints in MIDI can be up to 41, with the majority between 10 and 30 per image (Figure 5).
@@ 3.1 PHI Detection Pipeline
The general design of the PHI detection pipeline comprises three major modules: text localization, text extraction, and text analysis (see Figure 1). The text localization module can be an object detection model that identifies pixels containing text in the input image and returns bounding box coordinates. Once the text coordinates are determined, the text extraction module converts the text within these coordinates into machine-readable text, using OCR algorithms. The machine-encoded text is then analyzed semantically by a language model, which classifies whether it contains any PHI. In the last step, the pipeline flags the image as containing PHI, prompting further review and potential redaction.
Text localization
@@
YOLOv11 [21] is the latest version in the series of YOLO models [22], offering superior performance in real-time object detection tasks. The YOLOv11 architecture is a single-shot object detector consisting of three primary components: a backbone, a neck, and prediction heads. The backbone extracts image features, which are then processed and aggregated across multiple scales in the neck. The final predictions are produced by classification and detection heads. The main improvement in YOLOv11 over the previous version is an optimized backbone and neck, which enhance the extracted features and enable more precise object detection. For our study, we use the small variant of YOLOv11 pre-trained on the COCO dataset and fine-tune it on the RadPHI-train dataset. The model is trained to predict bounding boxes around a single line of text in the image, as illustrated in Figure 1. In addition, we combine real medical images with synthetic images of diverse background colors and shapes to help the model remain robust across all modalities. We use the ultralytics package [21] to train the model with a fixed image size of 640x640 for 100 epochs. The other training arguments are kept as default in the package. Apart from regular geometric augmentation techniques, the images undergo mosaic augmentation, where a new image is created based on patches of original images, as shown in Figure C1 (bottom right). The mosaic augmentation helps the model to generalize better to diverse backgrounds and object locations.
Text extraction
@@
EasyOCR [23] is an open-source OCR package supporting over 80 languages. It comprises two primary modules: recognition and detection. The detection module, analogous to our localization module, identifies text regions in an image using the Character Region Awareness For Text detection (CRAFT) model [24]. The recognition module [25] extracts text by leveraging the Scene Text Recognition module, a multi-stage model that transforms input images, extracts features, aggregates contextual information, and finally outputs the characters. In this study, we use EasyOCR as an out-of-the-box tool without re-training or fine-tuning its components. We keep the values of all parameters as default except for the low-bound text threshold. We reduced this threshold to 0.2 to include more space around detected texts by the CRAFT model, which we found helpful in avoiding missing texts as single letters like markers (e.g., L or R) or abbreviations of gender (e.g., M or F or D). The input to EasyOCR can be either a full image or a cropped image of the text region detected by the localization model. In our experiments, we compare both variants, see Section 4.
Text analysis
@@
GPT-4o [14] is OpenAI’s flagship language model in the Generative Pre-trained Transformer (GPT) family. These models are pre-trained on extensive unlabeled text datasets to generate human-like responses. GPT models have demonstrated their versatility in information extraction and NER tasks, particularly in biomedical text analysis [11]. Leveraging the superior performance of GPT in information extraction, we use GPT-4o to analyze the context of extracted texts from images in order to determine whether they contain PHI and which type of PHI is present. The extracted texts from the text extraction module are provided to GPT-4o as a list. The system prompt consists of the definition of PHI and non-PHI texts and instructions to evaluate the aggregated context of the input text rather than analyzing each line in isolation. Additionally, we provide examples in the prompt to make the model aware of ambiguous cases, following the few-shot prompting technique [26]. For instance, Age: or 65 contains a placeholder for age or hints at a number which could be age, but neither is considered PHI. To ensure consistent outputs, we constrain GPT-4o to produce a dictionary of mandatory fields through OpenAI’s pydantic tool, as outlined in Table 2. If the PHI is detected, the type field specifies one of the eight PHI classes or other if the PHI category is not included in the provided list. If no PHI is detected, the text analyzer assigns non-phi to the type field. Additional fields include raw_text, reason, and language, which store the raw extracted text, the justification for classifying the text as PHI or non-PHI, and the detected language of the original text, respectively. The temperature parameter is set to 0 to yield more deterministic responses.
@@
type str One of [date, identifier, name, address, age, gender, phone_nr, email, other, non-phi]

### v4

@@ Abstract
De-identification of medical images is a critical step to ensure privacy during data sharing in research and clinical settings. The initial step in this process involves detecting Protected Health Information (PHI), which can be found in image metadata or imprinted within image pixels. Despite the importance of such systems, there has been limited evaluation of existing AI-based solutions, creating barriers to the development of reliable and robust tools. In this study, we present an AI-based pipeline for PHI detection, comprising three key modules: text detection, text extraction, and text analysis. We benchmark three models—YOLOv11, EasyOCR, and GPT-4o— across different setups corresponding to these modules, evaluating their performance on two different datasets encompassing multiple imaging modalities and PHI categories. Our findings indicate that the optimal setup involves utilizing dedicated vision and language models for each module, which achieves a commendable balance in performance, latency, and cost associated with the usage of Large Language Models (LLMs). Additionally, we show that the application of LLMs not only involves identifying PHI content but also enhances OCR tasks and facilitates an end-to-end PHI detection pipeline, showcasing promising outcomes through our analysis.
@@ 1 Introduction
Protected Health Information (PHI) is health-related data that can identify an individual, often found in medical records, images, or any documents and data elements generated during participation in healthcare services. Before sharing data, e.g., for research purposes, it is crucial to de-identify or pseudonymize it to minimize the risk of re-identification. This de-identification step is mandatory for studies to comply with the Health Insurance Portability and Accountability Act (HIPAA) [1] and the General Data Protection Regulation (GDPR) [2], and falls under the responsibility of the data processor. Furthermore, fully de-identified datasets enable broader and unrestricted use, facilitating collaborative research and benchmarking without the need for complex and time-consuming processes to obtain patient authorization and approval.
De-identification is a complex process that involves many data types and formats, typically including a detection step followed by the redaction of the identified PHI. Medical images are a rich source of PHI, which can be embedded as metadata (e.g., in DICOM headers) or burned-in text within pixel data. However, manually detecting PHI in medical images is infeasible for large datasets, especially those that are curated for machine learning applications. Thus, the development of an automated tool for recognizing PHI is not just advantageous but essential for maintaining compliance and ensuring the privacy and security of sensitive health information. By leveraging automation, researchers and healthcare professionals can streamline the de-identification process, safeguarding patient privacy while enabling valuable insights from healthcare data.
@@
Nevertheless, the automatic detection of pixel-level PHI is challenging due to the diversity in imprint content, locations, font sizes, color schemes, annotations, and machine overlays. As the overall volume of medical data increases, so does the risk of incomplete PHI redaction, potentially leading to data breaches, which imply significant negative consequences for affected individuals and the healthcare sector, as well as high costs for responsible entities [3]. The importance of reliable PHI detection at scale and automated image quality control in general is strongly emphasized by the latest medical image data platform development program launched by the Advanced Research Projects Agency for Health (ARPA-H) [4].
@@
Previous works have employed Optical Character Recognition (OCR) technology to recognize text in medical images, leading to two primary approaches for handling burned-in text. The first approach leverages text detection or OCR models such as Tesseract [5] to localize burned-in texts and removes all of them, regardless of content [6, 7, 8]. While this method effectively eliminates any potential PHI, it also discards valuable information, such as measurements and markers. The second approach applies OCR models to extract burned-in texts and designs heuristic rules to recognize PHI based on matching the extracted text to a defined PHI dictionary and subsequently redact them [9, 10, 11]. While effective in structured settings, rule-based approaches are often tailored to specific imprint formats or language patterns and may struggle to generalize when faced with heterogeneous texts, ambiguous phrasing, or languages not covered by predefined rules. In addition, fine-grained classification of PHI is often preferred to coarse classification, which categorizes information simply as names or dates. This is particularly needed in certain research contexts where information such as visit dates or study-specific identifiers may be retained under a data use agreement, as it is essential for study analysis. As the volume and diversity of medical data continue to grow, advanced automatic solutions for PHI detection are essential, particularly for managing large datasets collected retrospectively. Therefore, a robust PHI detection tool must incorporate mechanisms that enable easy specification and adjustment of rules to meet the needs of different studies or country-specific guidelines.
@@
Our study focuses on AI-based approaches for automatically detecting PHI as burned-in pixels. Current commercial tools that address this need include MD.ai [12], Google [13], John Snow Labs [14], and Glendor [15]. Additionally, Microsoft provides an open-source tool called Presidio [16] that serves similar purposes. These tools commonly integrate two key components: a vision component that detects and converts burned-in texts into machine-encoded strings, and a language component that identifies any PHI content in the extracted texts [17]. To achieve this, OCR models are utilized for text recognition, while name-entity recognition (NER) models are employed to classify the PHI content. Recently, large language models (LLMs) have become the preferred candidates for the language component due to their flexible configurability and ability to understand complex language context in NER tasks, which has proven helpful in detecting PHI in biomedical text [17]. Despite some general remarks concerning the deployment of these tools [18, 19], particularly on the non-deterministic behavior and cost of LLMs, there is a lack of comprehensive studies dissecting the setup of vision and language models within a PHI detection pipeline. In addition, as multimodal LLMs with vision capabilities are being released today [20], the question remains whether they can be used to recognize PHI through end-to-end processing of images. To address the above concerns, we make the following contributions in our study:
@@
  • By leveraging publicly available medical imaging data and the HIPAA guidelines, we create two benchmark datasets comprising 1000 and 200 images, respectively. These datasets encompass various radiology modalities, including X-Ray, CT, PET, MRI, bone scintigraphy, mammography, and ultrasound. For both datasets we provide detailed annotations to support development and benchmarking within the research community.
  • Based on an extensive quantitative analysis, we provide insights into the optimal pipeline setup, considering performance, latency, and inference cost in terms of the number of generated tokens. Specifically, we explore the full capabilities of GPT-4o for both vision and language tasks, and analyze the trade-offs involved in utilizing it partially or entirely in the PHI detection pipeline.
@@ 2 Datasets
For the study at hand, we generate and utilize three distinct datasets: RadPHI-train as our training dataset, and two evaluation datasets, RadPHI-test and MIDI. RadPHI-train and RadPHI-test are created by overlaying synthetically generated imprints on curated medical images across various modalities. MIDI is created by overlaying information stored in DICOM tags, which may contain PHI, utilizing an industry standard DICOM viewer. The RadPHI-train is used to develop our text detection model, while the RadPHI-test is employed to evaluate different configurations of the PHI detection pipeline. Additionally, the MIDI dataset is reserved as a hold-out test set to assess the optimal configuration identified through the RadPHI-test evaluation. In the upcoming sections, we will elaborate on the generation process for each of these datasets.
@@ 2.1 PHI definition
For data simulation and evaluation, we define a list of PHI and non-PHI items based on the HIPAA guidelines [1] as detailed in Table 1. The six PHI categories include date, general identifier, patient name, address, phone number, and email. We group all identification numbers into a group of identifiers, including patient ID, insurance number, social security number, and any number or series that can be used to identify an individual. The ten non-PHI categories include age < 90, gender, height, weight, examination type, hospital, marker, scanner, diagnosis, and imaging personnel.
@@
Date All elements of dates related to an individual, including birth date, admission date, discharge date, and all ages over 89 years, as well as all elements of dates indicative of such age PHI DOB 01-01-2023
Identifier All identifiers related to an individual, such as patient identifier, insurance number, and medical record number PHI Patient ID: 0000.0001
@@
Age < 90 Age that is below 90, as well as all elements of dates indicative of such age. Non-PHI Age: 60
Gender The gender of an individual, e.g., male, female, or diverse, or the abbreviations such as m/f/d Non-PHI [M]
@@
Table 1: Imprint categories and examples. There are 16 categories, six of which are classified as PHI.
@@ 2.3 RadPHI-train
The main use of the RadPHI-train dataset is to enhance the robustness of the text localization module (see Section 3) against diverse radiological image backgrounds, varying numbers of imprints, and different imprint representations. This dataset is generated based on images from TotalSegmentator, BS-80K, and ChestX-ray8, applying the preprocessing techniques outlined in Section 2.2. Additionally, we simulate imprints on synthetic backgrounds, incorporating plain grayscale colors along with or without multiple boxes in various shades. The text localization module is trained to manage diverse representations of the imprint rather than focusing on its text content. To achieve this, we continuously simulate the imprints using different fonts, sizes, colors, and locations. These imprints may overlap with each other or with anatomical structures. Although such overlaps are not typically seen in real-world data, we intentionally include them as challenging cases for the model to adapt to. A sample from this dataset is shown in Figure C1 (left) in the Supplementary Material. In total, we generate 6,000 images, which are then divided into training and validation sets using an 80%-20% split.
@@ 2.4 RadPHI-test
Figure 5: Distribution of PHI and non-PHI imprints in RadPHI-test and MIDI datasets. In RadPHI-test, the number of imprints is limited to eight. In contrast, MIDI has significantly more imprints, correlating to the various DICOM tags and modality-specific imprints randomly displayed on the viewer.
@@ 2.5 MIDI
The MIDI dataset is designed to be a challenging hold-out dataset with realistic imprint visualization and placement via a modern DICOM viewer rendering. The image data is curated from the validation set of the 2024 Medical Image De-Identification Benchmark (MIDI-B) challenge [25]. This dataset originally consists of 228 DICOM images across multiple modalities, each containing synthetic PHI content embedded at both the DICOM header and pixel level. We utilize a DICOM viewer, specifically MD.ai [26], to overlay the DICOM tags onto the images. We randomly sample DICOM tags to ensure that the generated imprints represent all possible PHI categories, similar to the RadPHI-test dataset. After applying the overlays, we export the images from the viewer. The resulting images may include not only the DICOM tag overlays but also burn-ins by the challenge organizers. The final version of the dataset comprises 200 images categorized into five PHI types. The category Email is omitted because emails are neither burned in nor inserted into the header of the original dataset. In contrast to the RadPHI-test dataset, each image can have multiple imprints of the same PHI category. For instance, as illustrated in Figure 6, dates may appear in various contexts, such as the study date, image series, or even the patient comments. We performed instance-level annotation of the images by generating coordinates for PHI instances along with their corresponding categories. This annotation process was carried out and validated by two independent annotators to ensure accuracy and reliability. Figure 4 displays the distribution of PHI categories at the image level. Dates and patient names are among the most frequently occurring PHI elements in the dataset. While we intentionally limit the number of imprints to eight in RadPHI-test, the number of imprints in MIDI can be up to 41, with the majority between 10 and 30 per image (Figure 5).
@@ 3 Methods
The general design of the PHI detection pipeline comprises three major modules: text localization, text extraction, and text analysis (see Figure 1). The text localization module can be an object detection model that identifies pixels containing text in the input image and returns bounding box coordinates. Once the text coordinates are determined, the text extraction module converts the text within these coordinates into machine-readable text, using OCR algorithms. The machine-encoded text is then analyzed semantically by a language model, which classifies whether it contains any PHI. In the last step, the pipeline flags the image as containing PHI, prompting further review and potential redaction. In our study, we explore three distinct models: YOLOv11, EasyOCR, and GPT-4o. These models are configured in various combinations to fulfill the roles of the three modules mentioned above. We will describe each model in detail and outline how they are integrated to perform the functions of text localization, text extraction, and text analysis.
YOLOv11
@@
YOLOv11 [27] is the latest version in the series of YOLO models [28], offering superior performance in real-time object detection tasks. In our study, we employ YOLOv11 exclusively for the text localization module. The YOLOv11 architecture is a single-shot object detector consisting of three primary components: a backbone, a neck, and prediction heads. The backbone extracts image features, which are then processed and aggregated across multiple scales in the neck. The final predictions are produced by classification and detection heads. The main improvement in YOLOv11 over the previous version is an optimized backbone and neck, which enhance the extracted features and enable more precise object detection. For our study, we use the small variant of YOLOv11 pre-trained on the COCO dataset and fine-tune it on the RadPHI-train dataset. The model is trained to predict bounding boxes around a single line of text in the image, as illustrated in Figure 1. In addition, we combine real medical images with synthetic images of diverse background colors and shapes to help the model remain robust across all modalities. We use the ultralytics package [27] to fine-tune the model with a fixed image size of 640x640 for 100 epochs. The other training hyperparameters are kept as default in the package. Apart from regular geometric augmentation techniques, the images undergo mosaic augmentation, where a new image is created based on patches of original images, as shown in Figure C1 (bottom right). The mosaic augmentation helps the model to generalize better to diverse backgrounds and object locations.
EasyOCR
@@
EasyOCR [29] is an open-source OCR package supporting over 80 languages. It comprises two primary modules: recognition and detection. The detection module, analogous to our localization module, identifies text regions in an image using the Character Region Awareness For Text detection (CRAFT) model [30]. The recognition module [31] extracts text by leveraging the Scene Text Recognition module, a multi-stage model that transforms input images, extracts features, aggregates contextual information, and finally outputs the characters. We investigate EasyOCR under two distinct setups: (1) as a dedicated text extractor, and (2) as both a text localizer and a text extractor. For both setups, we use EasyOCR as an out-of-the-box tool without re-training or fine-tuning its components. We keep the values of all parameters as default except for the low-bound text threshold. We reduced this threshold to 0.2 to include more space around detected texts by the CRAFT model, which we found helpful in avoiding missing texts as single letters like markers (e.g., L or R) or abbreviations of gender (e.g., M or F or D). The input to EasyOCR can be either a full image or a cropped image of the text region detected by the localization model, depending on the experiment (see Section 4).
GPT-4o
@@
GPT-4o [20] is OpenAI’s flagship language model in the Generative Pre-trained Transformer (GPT) family. These models are pre-trained on extensive unlabeled text datasets to generate human-like responses. GPT models have demonstrated their versatility in information extraction and NER tasks, particularly in biomedical text analysis [17]. In our experiments, GPT-4o is configured to perform three distinct tasks: (1) text analysis, (2) text extraction, and (3) end-to-end image analysis. In the text analysis configuration, GPT-4o examines the context of extracted texts from images in order to determine whether they contain PHI and which type of PHI is present. The extracted texts from the text extraction module are provided to GPT-4o as a list. The system prompt consists of the definition of PHI and non-PHI texts and instructions to evaluate the aggregated context of the input text rather than analyzing each line in isolation. When GPT-4o acts as a text extractor, the prompts are structured to ensure that it returns the text visible in each cropped region predicted by YOLOv11. In the end-to-end image analysis scenario, GPT-4o is presented with a full image, prompted to perform text recognition, and tasked with analyzing each imprint identified in the image. It’s important to know that GPT-4o is not a dedicated object detection model, it does not return coordinates for the imprint. For the analysis task, we provide examples in the prompt to make the model aware of ambiguous cases, following the few-shot prompting technique [32]. For instance, Age: or 65 refers to a placeholder for age or hints at a number which could be age, but neither is considered PHI. To ensure consistent outputs, we constrain GPT-4o to produce a dictionary containing mandatory fields through OpenAI’s pydantic tool, as outlined in Table 2. If the PHI is detected, the type field specifies one of the eight PHI classes or other if the PHI category is not included in the provided list. If no PHI is detected, the text analyzer assigns non-phi to the type field. Additional fields include raw_text, reason, and language, which store the raw extracted text, the justification for classifying the text as PHI or non-PHI, and the detected language of the original text, respectively. The temperature parameter is set to 0 to yield more deterministic responses.
@@
type str One of [date, identifier, patient_name, address, phone_nr, email, other, non-phi]